the average age of 4 brothers is 12 .if the average age of their mother included the average is increased by 5 years.calculate the age of their mother....
Answer by stanbon(75887)

the average age of 4 brothers is 12 .if the average age of their mother included the average is increased by 5 years.calculate the age of their mother....
------
average = (sum of numbers)/(# of numbers)
-------
Equations:
sum of brothers ages/4 = 12
sum of brothers ages = 48
------
(sum of brothers + mother)/5 = 12+5
(48 + M) = 5*17
48 + M = 85
Mother's age = 37 yrs
-------------
Cheers,
Stan H.
--------------